OpenStudy (anonymous):

Basic Algebra equation solving:

OpenStudy (anonymous):

2x - 5 + 3

OpenStudy (michele_laino):

what is -5+3=?

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

hey!

OpenStudy (anonymous):

-2

OpenStudy (misty1212):

there is nothing to "solve" here because you do not have an equal sign you can compute \(-5+3=-2\) if you like, but solving requires some sort of equality

OpenStudy (anonymous):

Oh wait sorry guys! Lol. Its actually : 2x - 5 = 3

OpenStudy (misty1212):

lol how did i guess?

OpenStudy (misty1212):

add 5, then divide by 2!

OpenStudy (anonymous):

Ok, can u draw it out for me?

OpenStudy (michele_laino):

yhen, please add at both sides 5

OpenStudy (misty1212):

you can do it in your head

OpenStudy (misty1212):

what is 3+5?

OpenStudy (anonymous):

8 lol

OpenStudy (michele_laino):

oop.ss then add to both sides 5

OpenStudy (misty1212):

what is \(8\div 2\)?

OpenStudy (anonymous):

4

OpenStudy (misty1212):

you win!!

OpenStudy (michele_laino):

so what is your solution?

OpenStudy (anonymous):

yay! but how do u know what to add or subtract to both sides?

OpenStudy (misty1212):

\[2x-5=3\\ 2x-5+5=3+5\\ 2x=8\\ \frac{2x}{2}=\frac{8}{2}\\ x=4\]

OpenStudy (anonymous):

x = 4

OpenStudy (misty1212):

you know to add 5 because you want to get x by itself, so you have to get rid of that \(-5\) on the left

OpenStudy (anonymous):

Got it! omg u guys are so helpful!!! <3

OpenStudy (misty1212):

then you get \[2x=8\] and since \(2\times 4=8\) your answer is \(x=4\)

OpenStudy (michele_laino):

for example if you ahve this equation: x+ 5 = -8 then if I subtract 5 from both sides, I get: x= -13 then the last equation, is equivalent to the original equation. Please keep in mind that two equations are equivalent if and only f they admit the same solutions

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]

OpenStudy (anonymous):

Ok. I have another problem. Can u tell me if I get the right answer?

OpenStudy (misty1212):

kk

OpenStudy (anonymous):

Ok. 2x = 8

OpenStudy (michele_laino):

please divide both sides by 2

OpenStudy (anonymous):

x = 4 again!!

OpenStudy (misty1212):

amazing!

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

Yay! Ok i have another one.

OpenStudy (anonymous):

x = 4

OpenStudy (michele_laino):

it is already your solution

OpenStudy (anonymous):

Ok haha that was easy irk

OpenStudy (anonymous):

I'm posting another question. Will you help me with that one?

OpenStudy (michele_laino):

ok!